OpenStudy (anonymous):

For the graphs on the same axes, y=x, y=x^2, y=x^3, y=x^4, y=x^5, what is the growth pattern and what point of intersection do they have in common? Is it, aside from y=x, which is kinda in the middle, they are expanding? and they intersect at (0,0) ?

OpenStudy (anonymous):

this is what i got from google... wouldn't graph y=x for some reason haha but that would just be diagonally through the (0,0).. so yeah hehe :)

OpenStudy (usukidoll):

hmmm the graph expands.... when the power is high . y = x is a line

OpenStudy (anonymous):

okay yeah :) so would you just say that as the power rises, the functions expand, but y=x is just a line ?

OpenStudy (agent0smith):

x^n = 1 when x=1, for all n. x^n = 0 when x=0, for all n except 0

OpenStudy (agent0smith):

^^^so "what points of intersection" (1, 1) and (0, 0)

OpenStudy (anonymous):

ohh haha i see that now :P same as the last one! hahaha i didn't even see the (1,1) !

OpenStudy (anonymous):

okay so would these be correct? as the power gets bigger, the functions expand (with the exception of y=x, which is a line) and they intersect at point (0,0) and (1,1) ?

OpenStudy (anonymous):

@agent0smith is that right? :)

OpenStudy (agent0smith):

well they're all increasing, even y=x keeps rising Bigger powers just rise faster

OpenStudy (anonymous):

ohh okay awesome!!! thank you!!! :)

OpenStudy (triciaal):

y = (x)(x^n-1) where n is the next graph like a geometric sequence

OpenStudy (agent0smith):

^ that'd simplify to y=x^n